OpenStudy (ranya99):

If a rocket in space is moving at a constant velocity of 9.8 m/s and then uses its propulsion system to accelerate to 10. m/s during a 3.0 minute burn, what would be the acceleration of the rocket?

OpenStudy (irishboy123):

equation of motion \[v = u + at\] yields \[a = \dfrac{v-u}{t}\]

OpenStudy (anonymous):

1) the first thing you should know is that this rocket has a really bad burn rate. So if it starts with a velocity of 9.8 m/s and it accelerates to 10m/s in 3mins. The first thing we do is identify the problem, the rocket changes its speed by 0.2m/s and it takes 3mins to do that. All we have to do is divide the the change in speed by the time it took which is 3 mins or 180 seconds. and we get 0.2/180=0.0011111m/s^2. this seems like a small value so please double check the values you have given. 2) i do not have a lot of time to explain this is depth but we are going to use the equation s=ut+1/2at^2 you have given the values of displacement time and initial velocity, which respectively are 640m, 40s and 8m/s. All you need to do is put the values into the equation and then solve the equation for time.(t)